Atlas Inc. is a toy bicycle manufacturing company producing a five-inch small version of the bike that Lance Armstrong rode to win his first Tour de France. The assembly line at Atlas Inc. consists of seven work stations, each performing a single step. Stations and processing times are summarized here:

            • Step 1 (25 sec.): The plastic tube for the frame is cut to size.       

            • Step 2 (25 sec.): The tube is put together.                         

            • Step 3 (20 sec.): The frame is glued together                                 

            • Step 4 (30 sec.): The frame is cleaned.                               

            • Step 5 (35 sec.): Paint is sprayed onto the frame.                         

            • Step 6 (60 sec.): Wheels are assembled.                           

            • Step 7 (45 sec.): All other parts are assembled to the frame.                   

Under the current process layout, workers are allocated to the stations as shown here:  

            • Worker 1: Steps 1, 2           

            • Worker 2: Steps 3, 4                       

            • Worker 3: Step 5                             

            • Worker 4: Step 6                             

            • Worker 5: Step 7                             

Assume the workers are paid $10 per hour. Each bicycle is sold for $50 and includes parts that are sourced for $7.5. The company has fixed costs of $875 per hour. The demand rate is 40 per hour. Instructions: negative amounts should be indicated by a minus sign.

a. What is the cost of direct labor for the bicycle? Round your answer to the two decimals, for example, if your answer is 123.458, input 123.46.

Your answer is .

b. How much profit does the company make per hour? Round your answer to the nearest integer, for example, if your answer is 123.45, input 123.

Your answer is .

c. What would be the profits per hour if Atlas would be able to source the parts 10% cheaper? Round your answer to the nearest integer, for example, if your answer is 123.45, input 123.

Your answer is .

d. What would be the profits per hour if Atlas would be able to reduce fixed costs by 5%? Round your answer to the nearest integer, for example, if your answer is 123.45, input 123.

Your answer is .

e. What would be the profits per hour if Atlas would be able to reduce the processing time at the bottleneck by 5 seconds per unit? Round your answer to the nearest integer, for example, if your answer is 123.45, input 123.

Your answer is
Answer #1

a) Step-1

Calculate the capacity of each worker and the bottleneck capacity.

Worker-1 = 1/(25+25) = 1/50

Worker-2 = 1/(20+30) = 1/50

Worker-3 = 1/35

Worker-4 = 1/60

Worker-5 = 1/45

Bottlenecck the least capacity = Worker-4

Process capacity = Flow rate = 60 units / hr (1/60*3600)

Cost of direct labor= wages/time / flow rate/time

= [ ($10 * 5) / hr ] / [ 1 unit/min / 60 units/hr]= 50/60 = $0.83 / unit

b) Selling price = $50/unit

Parts Cost=$7.5/unit

Direct labor cost = $0.83/unit

Fixed cost = 875/hour

Demand = 40/ hour

So, total revenue in an hour = 50*40= $2000

Total cost in an hour = (7.5+0.83)*40+875= 1208.33

Total profit in an hour = 2000-1208.33=$791.67

c) If parts are 10% cheaper the parts cost will be = 7.5*0.9=6.75

The profit would be = 2000-((6.75+0.83)*40+875)= $821.67

d)Fixed cost reduced by 5%= 875*0.95= 831.25

Profit = 2000-((7.5+0.83)*40+831.25)= $835.42

e)

Bottleneck will be reduced by 5 seconds

Worker 4 will be = 1/55 = 3600*1/55= 65.45 units per hour

Direct labor cost = 10*5/65.45= 50/65.45= 0.76/unit

Profit = 2000-((7.5+0.76)*40+875)= $794.60

Please rate me

Thanks